OpenStudy (gawain):

Suppose the weights (w) of male students in MA102 are linearly related to their heights (h). If a student 62 inches tall weighs 120 pounds and a student 72 inches tall weighs 170 pounds, write an equation to express weight in terms of height.

jimthompson5910 (jim_thompson5910):

if we let x = height in inches y = weight in pounds then we'd have these two ordered pairs (62, 120) and (72, 170)

jimthompson5910 (jim_thompson5910):

(62,120) because x = 62 inches corresponds to y = 120 pounds (72,170) because x = 72 inches corresponds to y = 170 pounds

jimthompson5910 (jim_thompson5910):

Use the slope formula \[\Large m = \frac{y_2 - y_1}{x_2 - x_1}\] and tell me what slope you get

OpenStudy (gawain):

thanks

jimthompson5910 (jim_thompson5910):

what slope do you get?

OpenStudy (gawain):

m=5

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

y = mx+b y = 5*x + b ... replace m with 5 120 = 5*62 + b ... plug in (x,y) = (62,120) now solve for b and tell me what you get

OpenStudy (gawain):

b=-190

jimthompson5910 (jim_thompson5910):

correct, so because m = 5 and b = -190, this means the equation y =mx+b turns into y = 5x-190

OpenStudy (gawain):

thanks alot

jimthompson5910 (jim_thompson5910):

sure thing